OpenStudy (anonymous):

Can someone check my work? I have to calculate the surface area of a rectangular prism. The rectangular prism is 7in for length 1inch in height and 4in width I used the formula A = 2 (wh+lw+lh) and got 43inches

OpenStudy (anonymous):

Surface area = 2LW + 2(L+W)H. Try again.

OpenStudy (anonymous):

for a rectangular prism

OpenStudy (anonymous):

actually hold on

OpenStudy (anonymous):

Yea try again.

OpenStudy (anonymous):

But it's the same thing, don't worry

OpenStudy (anonymous):

hats the same thing

OpenStudy (anonymous):

Yes the two formulas are the same but you need to try it again because your answer is incorrect.

OpenStudy (anonymous):

What did I do wrong

OpenStudy (anonymous):

I'm not sure, but just use your formula and try it again. You will get a different answer:)

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

\[2((4 \times 1)+(7 \times 4)+ (7 \times 1))\] What does that give you?

OpenStudy (anonymous):

78?

OpenStudy (anonymous):

There you go:)

OpenStudy (anonymous):

Thank you I no what I did wrong

OpenStudy (anonymous):

Perfect.